# July 29, 2025

### v1

@@ 5 Discussion
For our trials with ControlNet we used a different implementation of our blending concept (6). Besides heuristics on semantic categories, we also used instance segmentation to calculate object size, and used this to retrieve small objects, which ControlNet tends to erase.
We also observed cases where sub-parts of objects were distorted, e.g. the tail lights of a vehicle. To monitor sub-parts we divided the ControlNet image into a grid of 16x32 cells (total 512) and measured the semantic distance to corresponding CARLA cells. The semantic distance was obtained with DreamSim [11], which uses embeddings of Vision Foundation Models to estimate a semantic distance between images. The distances are organized into a 16x32 thumbnail and resized back to the original image resolution. We used the inverse of the cells’ semantic distances as another quality map, after applying aggressive Gaussian smoothing to blur the grid boundaries ( 6). These three quality maps (semantic category, instance size, semantic distance) were all in the range [0, 1]. We combined them by multiplication, followed by dithering and gentle Gaussian smoothing to obtain a final quality map. This approach allowed us to obtain high fidelity images, but we decided to go with a simpler method since it was a bit too restrictive in regards to appearance variation, and the DreamSim calculations became a computation bottle neck.
In this context, we highlight another use case for DreamSim. ControlNet includes multiple guidance parameters, and we experimented with variants that introduce additional parameters, such as FreeU [35].
To automate Hyper-parameter Search (HPS), we selected a calibration set and computed the DreamSim distance between diffusion outputs and original images. Since our primary concern was hallucinations, and all outputs exhibited some degree of appearance variability, we optimized by minimizing the average DreamSim distance.
Figure 6: Alternative blending method. The final image (a) is obtained by a spatially varying weighted average of a CARLA image (b) and the ControlNet output (c). The output quality is estimated with DreamSim semantic distance (c), and heuristics on semseg (d) and instance size (e). These three quality maps are blended into a final map of blending weights (f), used as $w_{gen}$ in Eq. 1, Section 3.3 of the main paper.
@@
We observed that some adverse condition splits in ACDC yielded higher scores than evaluations on ACDC-Clear (see Section 4.1 and Table 1 in the main paper). To understand why object detection performed worse on clear images than on adverse images, we analyzed object distribution across different weather conditions. First, we depict the size distribution of objects in each weather condition in 8. Clear conditions contain the highest proportion of small objects, and the average object size is significantly smaller than in adverse conditions.
@@
This effect is further compounded by the fact that more difficult objects appear more frequently in certain weather conditions than others. To illustrate this, we analyze the object distribution across different weather conditions in ACDC (see 9).

### v2

@@ 5 Discussion
For our trials with ControlNet, we employed an alternative implementation of our blending concept. Further details will be disclosedin the future.
@@
We observed that some adverse condition splits in ACDC yielded higher scores than evaluations on ACDC-Clear (see Section 4.1 and Table 1 in the main paper). To understand why object detection performed worse on clear images than on adverse images, we analyzed object distribution across different weather conditions. First, we depict the size distribution of objects in each weather condition in 7. Clear conditions contain the highest proportion of small objects, and the average object size is significantly smaller than in adverse conditions.
@@
This effect is further compounded by the fact that more difficult objects appear more frequently in certain weather conditions than others. To illustrate this, we analyze the object distribution across different weather conditions in ACDC (see 8).